two particles A and B having opposite charges of 2*10^-6 and -2*10^-6 are placed at separation of 1 cm.calculate electric field at a point on axis of dipole 1 m away from the centre
 

Grade:12

1 Answers

Anshi
4 years ago
diapole moment p
p=q×d,
  =2×10-6×10-2
  =2×10-8
electric field at axis
  =2(k×p)÷r3
  =2(9×109×2×10-8)÷13
  =360